OpenStudy (anonymous):

Using complete sentences, explain how to factor each one. Be sure that the final factorization (or "answer") is a part of your explanation. 2x^2 + 13x + 15 2x^2 + 27x + 13

OpenStudy (dumbcow):

i'll show you how to do the 1st one 2x^2 +13x +15 -multiply the "x^2" coefficient and the constant....2*15 = 30 -Find 2 factors of 30 which add up to 13 .... 10+3 = 13 , 10*3 = 30 so i will use 10 and 3 , Note: if nothing works here then the expression cannot be factored -Next, rewrite expression replacing 13x with 10x+3x -group terms together (2x^2 +10x) +(3x +15) factor each group 2x(x+5) +3(x+5) factor out the (x+5) --> (x+5)(2x+3) done

OpenStudy (anonymous):

thank you

OpenStudy (dumbcow):

your welcome...did you figure out the other one

OpenStudy (anonymous):

not yet i was still working on the one you did trying to understand every little bit of it thanks you explaned it perfect and when i try the other one i will post it to let you look over it.